Fill in the missing blanks in the following table. Assume for simplicity that taxes are zero.

National Income and Real GDP​ (Y)

Consumption​ (C)

Saving​ (S)

​$12 comma 00012,000

​$8 comma 4008,400

​$nothing

​$13 comma 00013,000

​$9 comma 1009,100

​$nothing

​$14 comma 00014,000

​$9 comma 8009,800

​$nothing

​$15 comma 00015,000

​$10 comma 50010,500

​$nothing

​$16 comma 00016,000

​$11 comma 20011,200

​$nothing
Answer #1

Answer

The formula is:

S=Y-C

From the first blank

S=12000-8400=3600

S=13000-9100=3900 and so on

GDP C S
12000 8400 3600
13000 9100 3900
14000 9800 4200
15000 10500 4500
16000 11200 4800